OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (texaschic101):

first we find the slope by using the slope formula... slope(m) = (y2 - y1) / (x2 - x1) (-2,2)...x1 = -2 and y1 = 2 (5,1)...x2 = 5 and y2 = 1 now we sub slope(m) = (1 -2) / (5 - (-2) slope(m) = (-1) / (5 + 2) slope(m) = -1/7 now we use y = mx + b slope(m) = -1/7 use either of your points....(-2,2)....x = -2 and y = 2 now we sub 2 = -1/7(-2) + b 2 = 2/7 + b 2 - 2/7 = b 14/7 - 2/7 = b 12/7 = b so your equation is : y = -1/7x + 12/7.....but we need it in standard form Ax + By = C y = -1/7x + 12/7 --- add 1/7 to both sides 1/7x + y = 12/7 -- cannot have fractions, so we will multiply everything by 7 x + 7y = 12